# A classification of incompleteness statements

### v1

@@ Abstract.
Thanks to Hanul Jeon for discussion. The first author was partially supported by NSF Grant DMS-2054379.
@@ 1. Introduction
This latter theorem applies to $\Sigma^{1}_{1}$-definable theories—not just to the narrower class of $\Sigma^{0}_{1}$-definable theories—whence it is not an immediate consequence of Gödel’s original theorem.
@@ X Y Z
To place the ✗s on the table we show how to give appropriately non-standard definitions of arbitrarily strong sound theories. Theorem 1.2 places the first ✓ on the table; for this a “sufficiently strong” theory is any extension of $\Sigma^{1}_{1}\text{-}\mathsf{AC}_{0}$. To place the second ✓ on the table we forge a connection between provable $\Sigma^{1}_{1}$-soundness and a kind of “pseduo-ordinal analysis.” Whereas $\Pi^{1}_{1}$-soundness provably follows from the well-foundedness of a theory’s proof-theoretic ordinal, we show that $\Sigma^{1}_{1}$-soundness provably follows from the statement that a certain canonical ill-founded linear order lacks hyperarithmetic descending sequences. For the second ✓ a “sufficiently strong” theory is any extension of $\Sigma^{1}_{2}\text{-}\mathsf{AC}_{0}$.
@@ 2.2. Intermediate Cases
We can resolve two more cases with a subtler version of the proof of Theorem 2.1. First, we recall the following useful lemma. For $\Gamma\in\{\Sigma^{1}_{1},\Pi^{1}_{1}\}$, we let $\widehat{\Gamma}$ be the dual complexity class.
@@
$\tau(x)$ $:=x=U(x)\vee\Big{(}\mathsf{RFN}_{\widehat{\Gamma}}\big{(}U+\mathsf{RFN}_{\Gamma}(U)\big{)}\wedge\varphi(x)\Big{)}$
@@ 2.4. Hardest Case
No $\Sigma^{1}_{1}$-sound and $\Pi^{1}_{1}$-definable extension of $\Sigma^{1}_{2}\text{-}\mathsf{AC}_{0}$ proves its own $\Sigma^{1}_{1}$-soundness.
Theorem 1.2 was proved using concepts from ordinal analysis. In short, a connection is forged between $\Pi^{1}_{1}$-soundness and well-foundedness of proof-theoretic ordinals. Since we are now interested in $\Sigma^{1}_{1}$-soundness, we forge an analogous connection between $\Sigma^{1}_{1}$-soundness and pseudo-well-foundedness, where an order is pseudo-well-founded if it lacks hyperarithmetic descending sequences.
For the rest of this section assume that $T$ is a $\Sigma^{1}_{1}$-sound and $\Pi^{1}_{1}$-definable extension of $\Sigma^{1}_{2}\text{-}\mathsf{AC}_{0}$. In what follows, $\mathsf{PWF}(x)$ is a predicate stating that $x$ encodes a recursive pseudo-well-founded order (that is, a linear order with no hyperarithmetic decreasing sequence). Note that in $\mathsf{ATR}_{0}$, a subsystem of $\Sigma^{1}_{2}\text{-}\mathsf{AC}_{0}$, a universal quantifier over Hyp can be transformed into an existential set quantifier. This is [2, Theorem VIII.3.27] (originally proved by Spector and Gandy):
Theorem 2.8.
Let $\varphi(X)$ be a $\Sigma^{1}_{1}$ formula with no free set variables other than $X$. Then we can find an arithmetical formula $\theta(X,Z)$ such that
$\mathsf{ATR}_{0}\vdash\forall X\big{(}\varphi(X)\leftrightarrow\forall Z\leq_{H}X\;\theta(X,Z)\big{)}.$
Since $\Sigma^{1}_{2}\text{-}\mathsf{AC}_{0}$ contains $\mathsf{ATR}_{0}$, the statement $\mathsf{PWF}(x)$ is $T$-provably equivalent to a $\Sigma^{1}_{1}$ formula.
We will define $\prec_{T}$ to hold on pairs $(e,\alpha)$ where $e\in\textsc{Rec}$ and $\alpha\in dom(\prec_{e})$. We define $(e,\alpha)\prec_{T}(e^{\prime},\beta)$ to hold
if there is some $f\in\textsc{Hyp}$ so that $\mathsf{Emb}(f,\prec_{e}\upharpoonright\alpha+1,\prec_{e^{\prime}}\upharpoonright\beta)$ and $\mathsf{Pr}_{T}(\mathsf{PWF}(\prec_{e^{\prime}}))$.
@@
$\forall n\;\exists f\in\textsc{Hyp}\;\mathsf{Emb}\Big{(}(f,\prec_{e_{n}}\upharpoonright\alpha+1,\prec_{e_{n+1}}\upharpoonright\beta)\wedge\mathsf{Pr}_{T}\big{(}\mathsf{PWF}(\prec_{e_{n+1}})\big{)}\Big{)}.$
Let’s rewrite this:
$\forall n\;\exists f\Bigg{(}f\in\textsc{Hyp}\wedge\mathsf{Emb}\Big{(}(f,\prec_{e_{n}}\upharpoonright\alpha+1,\prec_{e_{n+1}}\upharpoonright\beta)\wedge\mathsf{Pr}_{T}\big{(}\mathsf{PWF}(\prec_{e_{n+1}})\big{)}\Big{)}\Bigg{.}$
Note that the predicate $f\in\textsc{Hyp}$ is $\Delta^{1}_{1}$ in $\mathsf{ACA}_{0}$ via the Kleene–Souslin Theorem [2, Theorem VIII.3.19]. Since $T$ is $\Pi^{1}_{1}$-definable, this means that the matrix of the formula above is generalized $\Pi^{1}_{1}$. Using choice for generalized $\Pi^{1}_{1}$ formulas (this is an application of $\Sigma^{1}_{2}\text{-}\mathsf{AC}_{0}$), we may swap the quantifiers in the above:
$\exists f\;\forall n\Big{(}f_{n}\in\textsc{Hyp}\wedge\mathsf{Emb}(f_{n},\prec_{e_{n}}\upharpoonright\alpha+1,\prec_{e_{n+1}}\upharpoonright\beta)\wedge\mathsf{Pr}_{T}\big{(}\mathsf{PWF}(\prec_{e_{n+1}})\big{)}\Big{)}.$
Note that the witness $f$ to this formula is technically a set that encodes countably many functions $(f_{n})_{n<\omega}$ in the standard way.
Using arithmetic comprehension, we form the composition $f_{\star}$ of the functions encoded in $f$. The function $f_{\star}$ is a hyp descending sequence in $\prec_{e_{0}}$. So $\prec_{e_{0}}$ is pseudo-ill-founded even though $T$ proves the opposite. Hence, $\neg\mathsf{RFN}_{\Sigma^{1}_{1}}(T)$. ∎
Before addressing the second claim, let’s record a dual form of Rathjen’s formalized version of $\Sigma^{1}_{1}$ bounding [1, Lemma 1.1].
@@
Since $\mathsf{PWF}$ (the set of pseduo-well-founded recursive linear orders) is $\Sigma^{1}_{1}$-complete, there is a total recursive function $\{k\}$ such that:
@@
By the recurision theorem and the S-m-n theorem, there is an integer $e$ so that $\mathsf{ATR}_{0}$ proves that $\forall i[\{e\}(i)\simeq\{\{k\}(e)\}(i)]$ (where $\simeq$ means that if either side converges then both sides converge and are equal). Working in $\mathsf{ATR}_{0}$, $\neg\mathsf{PWF}(e)$ implies $\neg\mathsf{PWF}(\{k\}(e))$, which implies $H(e)$, which implies $\mathsf{PWF}(e)$, which is a contradiction. So $\mathsf{ATR}_{0}\vdash\mathsf{PWF}(e)$. (Not that this implies $e\in\textsc{Rec}$ by the definition of $\mathsf{PWF}(e)$.)
@@
By Theorem 2.8 there exists a $\Pi^{1}_{1}$ formula $\pi(x)$ such that:

### v3

@@ Abstract.
Thanks to Hanul Jeon for discussion. Thanks to Mateusz Łełyk for discovering errors in an earlier draft. Thanks to an anonymous referee for helpful suggestions and, in particular, for suggesting the simple proof of Theorem 2.9, which we have included. The first author was partially supported by NSF Grant DMS-2054379.
@@ 1. Introduction
A theory is $\Pi^{0}_{1}$-sound (or, in general, $\Gamma$-sound) if all of its $\Pi^{0}_{1}$ theorems ($\Gamma$ theorems) are true. This notion can be formalized in the axiom systems we consider (see Definition 2.1).
@@
Note that this latter theorem applies to all $\Sigma^{1}_{1}$-definable theories and not just to the narrower class of $\Sigma^{0}_{1}$-definable theories.
@@ X Y Z
To place the ✗s on the table we show how to give appropriately non-standard definitions of arbitrarily strong sound theories. Theorem 1.2 places the first ✓ on the table; for this a “sufficiently strong” theory is any extension of $\Sigma^{1}_{1}\text{-}\mathsf{AC}_{0}$. For the second ✓ a “sufficiently strong” theory is any extension of $\mathsf{ATR}_{0}$.
[…]
@@ 2.1. Simplest Cases
Here $\mathsf{True}_{\Gamma}$ is a $\Gamma$-definable truth-predicate for $\Gamma$-formulas. For the complexity classes that we consider this truth-predicate is available already in the system $\mathsf{ACA}_{0}$.
For $\Gamma\in\{\Sigma^{1}_{1},\Pi^{1}_{1}\}$, we let $\widehat{\Gamma}$ be the dual complexity class. The following result is an immediate consequence of this definition:
Proposition 2.2.
Provably in $\mathsf{ACA}_{0}$, for $\Gamma\in\{\Sigma^{1}_{1},\Pi^{1}_{1}\}$, $T$ is $\Gamma$-sound if and only if $T+\varphi$ is consistent for every true $\widehat{\Gamma}$ sentence $\varphi$.
@@ 2.2. Intermediate Cases
We can resolve two more cases with a subtler version of the proof of Theorem 2.3. First, we recall the following useful lemma.
@@
$\tau(x)$ $:=U(x)\vee\Big{(}\mathsf{RFN}_{\widehat{\Gamma}}\big{(}U+\mathsf{RFN}_{\Gamma}(U)\big{)}\wedge\varphi(x)\Big{)}$
@@ 2.4. Hardest Case
No $\Sigma^{1}_{1}$-sound and $\Pi^{1}_{1}$-definable extension of $\mathsf{ATR}_{0}$ proves its own $\Sigma^{1}_{1}$-soundness.
First we give a short proof that was discovered by an anonymous referee:
Proof.
Let $T$ be a $\Sigma^{1}_{1}$-sound and $\Pi^{1}_{1}$-definable extension of $\mathsf{ATR}_{0}$ that proves its own $\Sigma^{1}_{1}$-soundness. Let $\Phi$ be the (conjunction of) the finitely many statements used in the proof (assume that a single sentence axiomatizing $\mathsf{ATR}_{0}$ is among them). The sentence $\Phi\in T$ is true $\Pi^{1}_{1}$. Hence, $\Phi+\Phi\in T$ is consistent and $\Phi+\Phi\in T\vdash\mathsf{RFN}_{\Sigma^{1}_{1}}(T)$. By running this same argument inside $\Phi+\Phi\in T$, we conclude that $\Phi+\Phi\in T\vdash\mathsf{Con}(\Phi+\Phi\in T)$. Yet $\Phi+\Phi\in T$ is a consistent and finitely axiomatized extension of $\mathsf{ATR}_{0}$, which contradicts Gödel’s second incompleteness theorem. ∎
Note that a dual version of this proof also establishes Theorem 1.2.
For the rest of this section we will give an alternate proof. In [4], Theorem 1.2 was proved using concepts from ordinal analysis. In short, a connection is forged between $\Pi^{1}_{1}$-soundness and well-foundedness of proof-theoretic ordinals. Since we are now interested in $\Sigma^{1}_{1}$-soundness, we forge an analogous connection between $\Sigma^{1}_{1}$-soundness and pseudo-well-foundedness, where an order is pseudo-well-founded if it lacks hyperarithmetic descending sequences.
For the rest of this section assume that $T$ is a $\Sigma^{1}_{1}$-sound and $\Pi^{1}_{1}$-definable extension of $\mathsf{ATR}_{0}$. In what follows, $\mathsf{PWF}(x)$ is a predicate stating that $x$ encodes a recursive pseudo-well-founded order (that is, a linear order with no hyperarithmetic decreasing sequence). A universal quantifier over Hyp can be transformed into an existential set quantifier in the theory $\mathsf{ATR}_{0}$. It follows that the statement $\mathsf{PWF}(x)$ is $T$-provably equivalent to a $\Sigma^{1}_{1}$ formula.
We will define $\prec_{T}$ to hold on pairs $(e,\alpha)$ where $e\in\textsc{Rec}$ and $\alpha\in dom(\prec_{e})$. We define $(e,\alpha)\prec_{T}(e^{\prime},\beta)$ to hold if
there is some $f\in\textsc{Hyp}$ so that $\mathsf{Emb}(f,\prec_{e}\upharpoonright\alpha+1,\prec_{e^{\prime}}\upharpoonright\beta)$ and $T\vdash\mathsf{PWF}(\prec_{e^{\prime}})$.
Here we write $\prec_{e}\upharpoonright\alpha+1$ for the restriction of the relation $\prec_{e}$ to $\{\gamma\in dom(\prec_{e})\mid\gamma\preceq_{e}\alpha\}$.
@@
$\forall n\;\exists g\in\textsc{Hyp}\;\mathsf{Emb}(g,f(n+1),f(n)).$
where we abuse notation to write $\mathsf{Emb}(g,f(n+1),f(n))$ for $\mathsf{Emb}(g,\prec_{e_{n+1}}\upharpoonright\beta_{n+1}+1,\prec_{e_{n}}\upharpoonright\beta_{n})$ to emphasize the role of $f$ in the statement.
By the Kleene–Souslin Theorem [3, Theorem VIII.3.19] in $\mathsf{ACA}_{0}$, $f$ is $\Delta^{1}_{1}$ since $f$ is in Hyp, so the formula $\mathsf{Emb}(g,f(n+1),f(n))$ is equivalent to a $\Sigma^{1}_{1}$ formula (this is an application of $\Sigma^{1}_{1}\text{-}\mathsf{AC}_{0}$, which is a consequence of $\mathsf{ATR}_{0}$ [3, Theorem V.8.3]).
$\mathsf{ATR}_{0}$ proves that Hyp satisfies $\Sigma^{1}_{1}$ choice, and therefore proves
$\exists g\in\textsc{Hyp}\;\forall n\;\mathsf{Emb}(g_{n},f(n+1),f(n)).$
Note that $g$ is technically a set encoding the graphs of the countably many functions $g_{n}$ in the usual way.
Using arithmetic comprehension, we form the composition $g_{\star}$ of the functions encoded in $g$$g_{\star}(0)=g_{0}(\beta_{1})$, $g_{\star}(1)=g_{0}(g_{1}(\beta_{2}))$ and so on. The function $g_{\star}$ is a hyp descending sequence in $\prec_{e_{0}}$, so $\prec_{e_{0}}$ is not pseudo-well-founded. Since $f(1)\prec_{T}f(0)$, we also have $T\vdash\mathsf{PWF}(\prec_{e_{0}})$. Recall that $\mathsf{PWF}(\prec_{e_{0}})$ is a $\Sigma^{1}_{1}$ claim. Hence, $\neg\mathsf{RFN}_{\Sigma^{1}_{1}}(T)$. ∎
Before addressing the second claim, let’s record a dual form of Rathjen’s formalized version of $\Sigma^{1}_{1}$ bounding [2, Lemma 1.1].
@@
[1, Theorem 1.3] implies that $\mathsf{PWF}$ (the set of pseduo-well-founded recursive linear orders) is $\Sigma^{1}_{1}$-complete (note that Harrison does not use self-reference or any other form of diagonalization in his proof). Hence, there is a total recursive function $\{k\}$ such that:
@@
By the recursion theorem and the S-m-n theorem, there is an integer $e$ so that $\mathsf{ATR}_{0}$ proves that $\forall i[\{e\}(i)\simeq\{\{k\}(e)\}(i)]$ (where $\simeq$ means that if either side converges then both sides converge and are equal). Working in $\mathsf{ATR}_{0}$, $\neg\mathsf{PWF}(e)$ implies $\neg\mathsf{PWF}(\{k\}(e))$, which implies $H(e)$, which implies $\mathsf{PWF}(e)$, which is a contradiction. So $\mathsf{ATR}_{0}\vdash\mathsf{PWF}(e)$. (Not that this implies $e\in\textsc{Rec}$ by the definition of $\mathsf{PWF}(e)$.)
@@
The formula $\exists f\in\textsc{Hyp}\;\mathsf{Emb}(f,\prec_{x},\prec_{T})$ consists of an existential hyp quantifier before a $\Pi^{1}_{1}$ matrix (the matrix is $\Pi^{1}_{1}$ since $\prec_{T}$ refers to provability in $T$ and $T$ is $\Pi^{1}_{1}$-definable). Hence, there exists a $\Pi^{1}_{1}$ formula $\pi(x)$ such that: